OpenStudy (anonymous):

show that (if 5n+3 is odd then n is even :n is a natural number)

OpenStudy (lgbasallote):

you can actually prove this using direct proof

OpenStudy (lgbasallote):

first, assume 5n + 3 is odd so 5n + 3 = 2k + 1 agree?

OpenStudy (lgbasallote):

http://www.youtube.com/watch?v=ILdBDOPoEDQ

OpenStudy (shubhamsrg):

well there'd be an alternative : 5n+3 = 2(2n+1) +n+1 now, 2(2n +1) is always even no matter what natural number n is only n+1 makes the difference clearly,in n+1 = 2k ,then whole no. is even and if n+1 = 2k +1 , then whole no. is odd we are looking for odd,, i.e. n+1 = 2k+1 or n = 2k i.e. n is even.. hope that was good enough,,